Please solve and show work.
Add the 2 matrices below, where possible.
3 -5
-5 + 1
-4 6
Thanks.

Answer by Fombitz(32388)

(3,-5,-4)+(-5,1,6)=(3-5,-5+1,-4+6)=(-2,-4,2)